# High-dimensional dynamics in low-dimensional networks

### v4

@@ Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive novel, precise conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
Introduction
@@
Recent work shows that many networks arising in nature and applications have low-dimensional structure [1]. This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
@@
In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15]. However, many of these recordings are made in the context of low-dimensional stimuli or behavior. A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20]. These observations raise the important point that the dimensionality of dynamics on a network likely depends on the structure of the network’s external input, in addition to the network’s internal structure.
@@
Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [1, 5, 12]. Other work considers external input, but assumes that the input is perfectly aligned with the network’s low-dimensional structure [3]. In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure. Other work [2, 4] considers the case of general external input, but the recurrent network is assumed to be weakly low-rank in the sense that the largest singular values are not asymptotically larger than the bulk of the singular values.
Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk. We show that such networks can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the detailed structure of the network connectivity. Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.”
We use mathematical analysis to derive precise conditions under which networks with low-rank structure produce high-dimensional dynamics and low-rank suppression. Our analysis shows that low-rank suppression is a general feature of networks with linearly stable dynamics and “strong” low-rank structure in the sense that the dominant singular values of the network are large.
On the other hand, high-dimensional dynamics are less general, and only arise when the network’s connectivity matrix has a non-vanishing component that satisfies the equal projector (EP) property. To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been pointed out in any previous work.
@@
We show that many common structural features such as biased weights, modularity, and spatial connectivity naturally produce connectivity matrices with the EP property, implying that these networks produce high-dimensional dynamics in addition to low-rank suppression. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
@@
Results
High-dimensional dynamics and low-rank suppression in a network with rank-one structure
For illustrative purposes, we begin with a simple, linear model (Figure 1a)
@@ Results
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
@@
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (bounded by $2\rho$; Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [25, 26].
Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [27]. If the dynamics were unstable, then $\|{\bf{z}}(t)\|$ would grow exponentially toward $\infty$, so we only consider parameter regimes with stable dynamics. Since strong low-rank structure also requires $|c|\gg\rho$, stability requires $c<0$ when $\rho=\mathcal{O}(1)$. In simulations here, we take $\rho=0.5$ and c=10c=-10. We consider more general network structures later.
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 10 in Appendix D). However, this conclusion does not necessarily carry over to recurrent networks.
@@
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The eigenvalues of $W$ have a dominant, negative term. d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}. g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
@@
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 10 in in Appendix D). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
Convergence to the steady-state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near $\bf{u}$}. Hence, the recurrent network suppresses inputs in the direction of $\bf{u}$}. A related explanation is that the Jacobian matrix J=WIJ=W-I of Eq. (1) has a large, negative eigenvalue near c1<0c-1<0 with associated eigenvector near $\bf{u}$} so the dynamics in Eq. (1) are highly compressive along $\bf{u}$}. This second explanation relies on the assumption that $W_{0}$ is a symmetric matrix (or at least a normal matrix) but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response. In the qausi-steady state,
@@
Since $W$ has a large singular value with left- and right singular values aligned to $\bf{u}$}, multiplication by $W$ amplifies the direction $\bf{u}$}. Therefore, the internal input, $W{\bf{z}}$, is large in the direction of $\bf{u}$} whenever $\bf{z}$} is moderate in the direction of $\bf{u}$}. In other words, $|{\bf{u}}\cdot W{\bf{z}}|\gg|{\bf{u}}\cdot{\bf{z}}|$. This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq. (3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}.
This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}. Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h). We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [21, 28] as we will show later.
Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
@@
Figure 2: Response properties of a non-normal recurrent network with rank-one structure. Same as Figure 1 except the low-rank part of $W$ is non-normal. The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c). The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
@@
In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example. This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$). While the variability along $\bf{v}$} itself was not much smaller than a random directions (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f). Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g). The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in a later section, but note that it is closely aligned to $\bf{u}$} whenever $c$ is large. In other words, inputs in the direction of ${\bf{x}}_{\textrm{aligned}}$, which is close to $\bf{u}$}, are suppressed by the network. Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. In the next section, we derive precise conditions for low-rank suppression and low-dimensional dynamics.
General conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple network with rank-one structure. Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and derive precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
Assumption 1: Dynamics are linear or near a stable equilibrium.

### v5

@@ Abstract
Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks would also be low-dimensional. Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input. However, networks in nature are rarely isolated. Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations. We find that dynamics in such networks can be high- or low-dimensional and we derive mathematical conditions on the network structure under which dynamics are high-dimensional. In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression. Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of its response to external input.
preprint: APS/123-QED
I Introduction
@@
Recent work shows that many networks arising in nature and applications have low-rank structure [1]. This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
@@
In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 2, 12]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15]. However, many of these recordings are made in the context of low-dimensional stimuli or behavior. A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20]. These observations raise the important point that the dimensionality of dynamics on a network depends on the structure of the network’s external input, in addition to the network’s internal structure.
@@
There is a large body of theoretical work on the dimensionality of dynamics on low-rank networks [21, 22, 2, 6, 3, 7, 5, 8, 9, 10, 1, 6, 23, 4, 24, 10, 9, 25, 11, 12, 2]. Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [2, 6, 1]. Other work considers external input, but assumes that the input is perfectly aligned with the network’s low-dimensional structure [4]. In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure. Other work [3, 5] considers the case of general external input to networks that are weakly low-rank in the sense that the largest singular values are $\mathcal{O}(1)$ in magnitude, consistent with connection weights that scale like $1/N$ where network size, $N$, is assumed to be large.
In this study, we consider networks with strongly low-rank structure in the sense that they have a small number of asymptotically large singular values, while the rest are small or moderate in magnitude. This is consistent with connection weights that are asymptotically larger than $1/N$ in magnitude, often taken to be $\mathcal{O}(1/\sqrt{N})$ in theoretical work. Importantly, we study the dimensionality of the dynamics on these networks when they are driven by external input. The dimensionality of activity driven by intrinsic dynamics is a fundamentally different question treated in previous work [1, 6, 2]. We also focus on dynamics near linearly stable equilibrium. We show by example that networks under these conditions can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the network’s structure. Perhaps counterintuitively, we also find that networks under these conditions can suppress inputs aligned to their low-dimensional structure relative to randomly aligned inputs, an effect we call “low-rank suppression.”
We use mathematical analysis to derive conditions under which low-rank suppression and high-dimensional dynamics can occur in strongly low-rank networks close to a stable equilibrium. Our analysis shows that low-rank suppression arises generally in these networks. High-dimensional responses to high-dimensional external inputs, on the other hand, arise when the dominant low-rank part of the network’s connectivity matrix has a non-vanishing component with the EP property (meaning that its column space is equal to its row space). Low-dimensional dynamics can arise in these networks through a form of non-normal amplification [26, 27, 28], but our mathematical analysis shows that this form of non-normal amplification can only occur when the network is strongly non-EP. Strong non-normality is not sufficient. To our knowledge, this connection between the EP property of matrices, non-normal amplification, and the dimensionality of recurrent network dynamics has not been described in previous work.
@@
We show that many common structural features such as biased weights, modularity, and spatial connectivity structure naturally produce strongly low-rank connectivity matrices with low-rank parts satisfying the EP property, implying that these networks produce low-rank suppression and high-dimensional responses to high-dimensional inputs. We also draw connections between low-rank suppression and the mathematical theory of balanced networks [29], extending previous work in this direction [4]. Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
@@
II High-dimensional dynamics and low-rank suppression in a network with rank-one structure
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The eigenvalues of $W$ have a dominant, negative term. d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}. g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). The norm of $\bf{z}$} is more than 11 times smaller in response to the aligned versus random perturbation. h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
For illustrative purposes, we begin with a simple, linear recurrent network model (Figure 1a)
@@ II High-dimensional dynamics and low-rank suppression in a network with rank-one structure
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ recurrent connectivity matrix, $W$, takes the form
@@
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (approximately bounded by $2\rho$; Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [33, 34].
Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [35]. If the dynamics were unstable, then $\|{\bf{z}}(t)\|$ would grow exponentially toward $\infty$, so we only consider parameter regimes with stable dynamics. Since strong low-rank structure also requires $|c|\gg\rho$, stability requires $c<0$ when $\rho=\mathcal{O}(1)$. In simulations here, we take $\rho=0.5$ and c=10c=-10. We consider more general models and network structures later.
We first provided a high-dimensional input, ${\bf{x}}(t)$. Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input to the network. Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}. Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 7 in Appendix F). However, this conclusion does not necessarily carry over to recurrent networks.
@@
which was generated identically to, but independently from $\bf{u}$}. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 7 in in Appendix F). In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
Convergence to the steady-state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near $\bf{u}$}. Hence, the recurrent network suppresses inputs in the direction of $\bf{u}$}. A related explanation is that the Jacobian matrix J=WIJ=W-I of Eq. (1) has a large, negative eigenvalue near c1<0c-1<0 with associated eigenvector near $\bf{u}$} so the dynamics in Eq. (1) are highly compressive along $\bf{u}$}. These explanations rely on the assumption that $W_{0}$ is a symmetric matrix, or at least a normal matrix, but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response. In the quasi-steady state, we have
@@
Since $W$ has a large singular value with left- and right singular values aligned to $\bf{u}$}, multiplication by $W$ amplifies the direction $\bf{u}$}. Therefore, the internal input, $W{\bf{z}}$, is large in the direction of $\bf{u}$} whenever $\bf{z}$} is moderate in the direction of $\bf{u}$}. In other words, $|{\bf{u}}\cdot W{\bf{z}}|\gg|{\bf{u}}\cdot{\bf{z}}|$. This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq. (3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}, as we saw in Figure 1e,g.
This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}. Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (i.e., ${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h). We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [29, 36] as we will show later.
III Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
Figure 2: Response properties of a non-normal recurrent network with rank-one structure. Same as Figure 1 except the low-rank part of $W$ is non-normal. The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c). The network exhibits low-dimensional dynamics (d), in contrast to Figure 1. The norm of $\bf{z}$} is more than 13 times smaller in response to the aligned versus random perturbation.
@@ III Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example. This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$). While the variability along $\bf{v}$} itself was not much smaller than a random direction (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f). Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g). The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in Section IV.5 and Appendix D, but note that ${\bf{x}}_{\textrm{aligned}}$ is closely aligned to $\bf{u}$} whenever $c$ is large. Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics. In the next section, we derive more general conditions for low-rank suppression and low-dimensional dynamics.
IV General conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple network with rank-one structure. Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions. We first state two assumptions that we make throughout this section. Then, we show that low-rank suppression occurs under these assumptions alone. Finally, we analyze the conditions for high-dimensional dynamics under various assumptions on ${\bf{x}}(t)$.
IV.1 Assumption 1: Dynamics are linear or near a stable equilibrium.